PartnerObserving Log for 2007 Jan 19/20

snowing whole day -- Main.fan - 20 Jan 2007

Observers: Rick Pogge & Xiaohui Fan
Inst. Support Astronomer: Dave Thompson
TOs: Aaron Ceranski & David Gonzalez

As of 1500MST, we're under a heavy snow warning through noon tomorrow. We'll exercise a few queue OBs tonight to get the system into our fingers. Outside is 10-15m/s winds and horizontal snow.

All times below are are UTC 2007-Jan-20

0315 - *mock observation of Clementini M31S2 B initiated

0325 - momentary power glitch during slew from horizon-pointing to target, switched briefly to UPS.

0357 - drives reset and restarted, but trouble turning on the enclosure rotation, J. Hill consulted, TOs working through it...

0625 - No luck getting the enclosure rotation back in operation (see IssueTrak #439) - we're officially done for the night even for mock observing.

-- RichardPogge - 20 Jan 2007
Topic revision: r12 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback